A telephone company charges $4.95 per month plus $0.05 per minute. The monthly cost of the
plan can be described by the equation 𝑐 = 0.05𝑚 + 4.95, where m is the number of minutes.
how to find y intercept with out graphing
If you talk 140 minutes in a month, how much will your plan cost

Answer by rothauserc(4718)

c = 0.05m + 4.95
:
The graph of c has m plotted on the x axis, therefore
:
we set m = 0 to find the y intercept
:
c = 0.05 * 0 + 4.95
:
***********************
the y intercept is 4.95
***********************
:
c = 0.05 * 140 + 4.95 = 11.95
:
*********************************************
The cost for 140 minutes in a month is $11.95
*********************************************
: